Pasoh Forest Reserve (10.01.2013 to 12.31.2013)

You are here

PDF versionPDF version
Year: 
2013
Quarter: 
October - December
Communications activities: 

None

Unusual Events at the TEAM Site: 

None

Protocol Activities: 

Vertebrate;

Cameras deployed in Array 3 were collected and images saved into project computer. Subsequently 18 cameras were set up in Array 1 (8 cameras were stolen while 4 had problems leaving only 18 cameras in working condition) in November 2013 and collected by the end of December 2013.

Climate:

In late October 2013, an updated program sent by the TEAM Technical team was installed into the datalogger to resolve the power shortage issue at the climate tower. The new program shuts off the fan after sunset (when temperatures are expected to be cooler) and starts again at dawn. Data collected for the month of November and December 2013 showed complete data logging and sufficient power to the climate station  throughout.

Protocol Problems: 

Vertebrate:

Pasoh FR area of 13,900ha can only fit in 57 camera locations with a 2km radius apart as per protocol. Thus some of the camera locations for the new Array 3 are located near forest edges and a few are even just outside of the reserve, but still under forest cover.

There are still issues (problems) with uploading of .tpk files (computer or system compatability, internet firewalls and security etc).

Climate:

The recent decision by the TEAM technical team to shut off the fan during dark hours was an effective saving measure and should be considered to be included in the protocol.

 

Schedule Problems: 

Vertebrate:

Due to the loss of cameras (theft and malfunction), sampling in Array 1 had to be conducted in 2 sessions, further delaying the sampling schedule by a month.

Due to problems with uploading .tpk files, we could not complete the vertebrate protocol within a desired time frame.

 

Logistical Problems: 

None

Suggestions: 

The recurring issues with uploading of .tpk files need to be resolved.